OpenStudy (lilai3):

ey need elp rite now pleaz... Of 8 households chores, in how many ways can you do three-fourths of them? Latanya volunteers to bake and deliver pastries to elderly people in her neighborhood. In how many different ways can Latanya deliver to 2 out of 6 elderly people in her neighborhood?

OpenStudy (campbell_st):

you need to do 6 of the 8 then its 8 x 7 x 6 x 5 x 4 x 3 =

OpenStudy (lilai3):

wait..@campbell_st, do you mean the first one?

OpenStudy (campbell_st):

yep the 1st one

OpenStudy (lilai3):

okay, thanks, then its continous, rite?

OpenStudy (lilai3):

8! rite?

OpenStudy (blacksteel):

This is very simple combinatorics. In the first problem, you want to do 3/4 of the 8 household chores, or 6 total. If the order you do them in matters, campbell's answer is correct. This is because under permutation, doing chore A first and B second would be a different case than doing chore B first and A second. Combination simply looks at the number of different sets regardless of order. To choose a items from b items total, the notation is a C b, and is given by the formula b!/(a!*[b-a]!) (where n! = n*n-1*n-2*n-3*...*3*2*1) For a = 6, b = 8, there are 8!/(6!*2!) = 8*7/2 = 28 ways to pick a set of 6 chores to do. Similarly, in the second problem if you consider visiting Mr. Red then Mr. Blue to be a different case than visiting Mr. Blue then Mr. Red, there are 6x5 = 30 possible ways to visit 2 people. If order doesn't matter, you use combination again: 6!/2!*4! = 6*5/2 = 15

OpenStudy (lilai3):

OMFG, LMFAO, thank you a billion tons..wow...it must've took hoooooourrrrssss.... but i very thank you as shown, and as said, for i also gave you medals and became your fan hahah and i still THANK YOU VERY BERRY MUCH!!!!!!!11 ;)

OpenStudy (anonymous):

Question 1: Answer is \[8!/(8-6)! = 20160\] if order of doing chores matters; \[8!/(8!)(8-6)! = 28\] if it doesn't. Question 2: Answer is \[6!/(6-2)!=30\] if order of delivery matters; \[6!/(2!)(6-2)!=15\] if it doesn't.

OpenStudy (lilai3):

thank you @rchstudy,also!!!!!! thank you!!!!! <333